# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ II Summary of Main Results
The second thermodynamic inequality we prove is the thermodynamic bound on the transport of any scalar observable $z(t)\equiv z(x_{\tau},\tau)$ [51]
$\frac{\langle z(t)-z(0)-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z(\tau)\rangle^{2}}{t\mathcal{D}^{\mathrm{(c)}}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}}{2}\,$, (16)
where $\mathcal{D}^{\mathrm{(c)}}_{t}$ is the so-called fluctuation-scale function accounting for how much the observation stretches microscopic coordinates. In continuous space it was shown [113, 51] that
$\mathcal{D}_{t}^{\mathrm{c}}=\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\left\langle\left[\nabla_{\bm{x}}z_{\tau}\right]^{T}\bm{\gamma}^{-1}(\tau)\nabla_{\bm{x}}z_{\tau}\right\rangle.$ (17)
For MJP we in turn find
$\mathcal{D}_{t}\equiv\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}(z_{y}(\tau)-z_{x}(\tau))^{2}r_{xy}p_{x}(\tau).$ (18)
The fluctuation-scale function for both, continuous and discrete space dynamics, is operationally accessible by considering increments $\mathrm{d}z(\tau)=z(x_{\tau+\mathrm{d}\tau},\tau+\mathrm{d}\tau)-z(x_{\tau},\tau)$ along a trajectory and recognizing (as in continuous space [113, 51]) that
@@
The last family of thermodynamic bounds we prove are thermodynamic bounds on observable correlations (or, in short, “correlation bounds”) as a generalization of the bounds in Ref. [113]. It involves correlations of time-independent observables $V_{\tau}\equiv V(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}V_{k}$ and $z_{\tau}\equiv z(x_{\tau})=\sum_{k}\delta_{k,x_{\tau}}z_{k}$ and their time-average along a trajectory, e.g., $\overline{V}_{t}=t^{-1}\int_{0}^{t}V_{\tau}\mathrm{d}\tau$, resulting in a lower bound of the integrated entropy production rate $\dot{S}_{\mathrm{tot}}(\tau)$ weighted by a “pseudo variance” $\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau})=\sum_{x,y}p_{xy}^{\mathrm{ps}}(\tau)(z_{x}+z_{y}-2F(\tau))$ [114]
@@
where $F(\tau)$ may be any time-(in)dependent function. The tilted two-point probability $p_{xy}^{\mathrm{ps}}(\tau)$ requires knowledge about transition rates, hence the pseudo variance is not operationally accessible. However, it can be controlled by choosing bounded observables. In NESS, Eq. (20) reads
@@
The response of a single-time observable $O(t)$ of a continuous-space diffusion to a perturbation in the noise amplitude, which effectively is a temperature perturbation [135], $\bm{\sigma}\to\bm{\sigma}+\delta\tilde{\bm{\sigma}}$ was recently derived using stochastic calculus and reads [136]
@@
The result in Eq. (26) relates the response to an observable $O$ to a general small perturbation of the dynamics to correlation functions of the unperturbed dynamics (which, notably, for both MJP and diffusions are not necessarily operationally accessible). In the case of equilibrium dynamics, Eq. (26) reduces to the Fluctuation Dissipation Theorem (FDT) in Eq. (30). Therefore, Eq. (26) can in some sense be seen as a (mathematical) extension of the FDT to irreversible and transient systems. In Appendix A.8 we generalize the response in Eq. (26) to perturbations of pathwise observables of MJP, e.g.
$O(t)=\int_{s=0}^{s=t}\mathrm{Tr}[\bm{b}(s)^{T}\mathrm{d}\bm{\varepsilon}(s)],$
for some time-dependent hollow matrix $\bm{b}(s)$.
@@ III.1 Stochastic Equations of Motion
Let $(\bm{x}_{\tau})_{0\leq\tau\leq t}$ be a trajectory of $\bm{x}_{\tau}\in\mathds{R}^{d}$ evolving according to overdamped Langevin dynamics [61]
@@
Consider now a discrete state space $\mathcal{N}\subseteq\mathds{N}$ which, without loss of generality, we can enumerate as $\mathcal{N}=\{1,2,\dots,N\}$. Focusing for the time being on time-homogeneous systems (i.e., time-independent transition rates $r_{ij}$ between states $i$ and $j$) the state probability vector $p_{i}(\tau)$ evolves according to the master equation [14, 8]
@@
where the summation is only over $j$. In terms in the incoming and outgoing fluxes $J_{xy}(\tau)=r_{xy}p_{x}(\tau)$ we have as $\frac{\mathrm{d}}{\mathrm{d}\tau}p_{i}(\tau){\color[rgb]{0,0,1}=}\sum_{j\neq i}[J_{ij}(\tau)-J_{ji}(\tau)]$, and in vector notation $\frac{\mathrm{d}}{\mathrm{d}\tau}\bm{p}(\tau)=\bm{L}\bm{p}(\tau)$ with generator $\bm{L}$ having entries $L_{ij}=\delta_{ij}r_{ii}+(1-\delta_{ij})r_{ji}$ and exit rate $r_{ii}=-\sum_{j\neq i}r_{ij}$. The transition probability (or “propagator”) is therefore given by $P(i,\tau^{\prime}-\tau|j)=P(i,\tau^{\prime}|j,\tau)=\left(\exp{(\tau^{\prime}-\tau)\bm{L}}\right)_{ij}$.
A path is fully determined by the sequence of states $(x_{i})$ and transition times $t_{i}$, $(x_{\tau})_{0\leq\tau\leq t}=(x_{0},t_{0}=0;x_{1},t_{1};\dots)$. The probability of such a trajectory is then [6, 36, 95, 138, 139]
$\mathds{P}\left[(x_{\tau})_{0\leq\tau\leq t}\right]=p_{x_{0}}(0)\mathrm{e}^{\sum_{k}\tau_{k}(t)r_{kk}}\prod_{i,j}(r_{ij})^{n_{ij}(t)},$ (33)
@@
$\underbrace{\mathrm{d}n_{xy}(\tau)}_{\text{``displacement"}}=\underbrace{r_{xy}\mathds{1}_{x_{\tau}=x}}_{\text{``drift"}}\mathrm{d}\tau+\underbrace{\mathrm{d}\varepsilon_{xy}(\tau)}_{\text{``noise"}}.$ (38)
Defining the matrix $(\bm{R}(x_{\tau}))_{ij}\equiv(1-\delta_{ij})r_{ij}\mathds{1}_{x_{\tau}=i}$, we can write Eq. (38) as a stochastic matrix equation
@@
Note that $\mathrm{d}\bm{\varepsilon}(\tau)$ is not Gaussian noise, but rather a shifted (due to the $\bm{R}(x_{\tau})\mathrm{d}\tau$ term) time-inhomogeneous Poisson process. Explicitly, because for $x_{\tau}=x$ and $y\neq x$ the differential $\mathrm{d}n_{xy}(\tau)\sim\mathrm{Poi}(r_{xy}\mathrm{d}\tau)$ is Poisson, the values $\mathrm{d}\varepsilon_{xy}(\tau)$ can take for any $x_{\tau}$ are $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau,2-r_{xy}\mathrm{d}\tau,3-r_{xy}\mathrm{d}\tau,\dots\}$, where $\mathrm{d}\varepsilon_{xy}(\tau)=0$ happens if $x_{\tau}\neq x$. In case $x_{\tau}=x$, one can identify $P(\mathrm{d}\varepsilon_{xy}(\tau)=k-r_{xy}\mathrm{d}\tau|x_{\tau}=x)=\mathrm{poi}_{r_{xy}\mathrm{d}\tau}(k)$ for any $k\in\mathds{N}_{\geq 0}$. In the limit of small $\mathrm{d}\tau$ to linear order in $\mathrm{d}\tau$ the noise therefore takes values $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau\}$ with probabilities $1-p_{x}(\tau),\left(1-r_{xy}\mathrm{d}\tau\right)p_{x}(\tau),r_{xy}\mathrm{d}\tau p_{x}(\tau)$, respectively. The probability for any $k\geq 2$ are $\mathcal{O}(\mathrm{d}\tau^{k})$ and can therefore be neglected.
@@
$=\delta_{xk}\mathds{1}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau).$
@@
The above correlators allow, for example, to evaluate the covariance between the number of jumps ate different space-time points along the trajectory, i.e., for $\tau\leq\tau^{\prime}$, [60]
@@
$\times\left[\delta(\tau-\tau^{\prime})\delta_{ik}\delta_{jl}+\mathds{1}_{\tau=\tau^{\prime}}\delta_{ik}r_{ij}+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}P(i,\tau^{\prime}|l,\tau)\right].$
@@
$\mathds{P}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}\mathrm{d}n_{xy}(s)\right],$ (44)
@@
Meaning Continuous Discrete
Displacement $\mathrm{d}\bm{x}_{\tau}$ $\mathrm{d}\bm{n}(\tau)$
Noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ $\mathrm{d}\bm{\varepsilon}(\tau)$

### v2

@@ II Summary of Main Results
The second thermodynamic inequality we prove is the thermodynamic bound on the transport of any scalar observable $z_{{\color[rgb]{0,0,0}\tau}}\equiv z(x_{\tau},\tau)$ [40]
$\frac{\langle z_{{\color[rgb]{0,0,0}t}}-z_{{\color[rgb]{0,0,0}0}}-\int_{0}^{t}\mathrm{d}\tau\partial_{\tau}z_{{\color[rgb]{0,0,0}\tau}}\rangle^{2}}{t\mathcal{D}^{\mathrm{(c)}}_{t}}\leq\frac{\Delta S_{\mathrm{tot}}^{\mathrm{(c)}}}{2}\,$, (18)
where in continuous space it was shown [30, 40] that
$\mathcal{D}_{t}^{\mathrm{c}}=\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\left\langle\left[\nabla_{\bm{x}}z_{\tau}\right]^{T}\bm{\gamma}^{-1}(\tau)\nabla_{\bm{x}}z_{\tau}\right\rangle,$ (19)
while for MJP we find
$\mathcal{D}_{t}\equiv\frac{1}{t}\int_{0}^{t}\mathrm{d}\tau\sum_{x,y}[z_{y}(\tau)-z_{x}(\tau)]^{2}r_{xy}p_{x}(\tau).$ (20)
Here, $\mathcal{D}^{\mathrm{(c)}}_{t}$ is the so-called fluctuation-scale function accounting for how much the observation stretches microscopic coordinates. For example, note that stretching $z_{\tau}\mapsto az_{\tau}$ for some constant $a\in\mathbb{R}$ does not alter Eq. (18) as the stretch in the nominator is canceled by the change in $\mathcal{D}^{\mathrm{(c)}}_{t}$. In Eq. (20), $z_{x}(\tau)$ is the specific value of the function $z_{\tau}$ if $x_{\tau}=x$, i.e., $z_{\tau}\equiv z(x_{\tau},\tau)=\sum_{k}\delta_{x_{\tau}k}z_{k}(\tau)$. The fluctuation-scale function for both, continuous and discrete space dynamics, is operationally accessible by considering increments $\mathrm{d}z(\tau)=z(x_{\tau+\mathrm{d}\tau},\tau+\mathrm{d}\tau)-z(x_{\tau},\tau)$ along a trajectory and recognizing (as in continuous space [30, 40]) that
@@
The last family of thermodynamic bounds we prove are thermodynamic bounds on observable correlations (or, in short, “correlation bounds”) as a generalization of the bounds in Ref. [30]. It involves correlations of time-independent observables $V_{\tau}\equiv V(x_{\tau})=\sum_{k}{\color[rgb]{0,0,0}\delta_{x_{\tau}k}}V_{k}$ and $z_{\tau}\equiv z(x_{\tau})=\sum_{k}{\color[rgb]{0,0,0}\delta_{x_{\tau}k}}z_{k}$ and their time-average along a trajectory, e.g., $\overline{V}_{t}=t^{-1}\int_{0}^{t}V_{\tau}\mathrm{d}\tau$, resulting in a lower bound of the integrated entropy production rate $\dot{S}_{\mathrm{tot}}(\tau)$ weighted by a “pseudo variance” $\mathrm{pvar}_{\mathrm{ps}}^{F}(z_{\tau})=\sum_{x,y}p_{xy}^{\mathrm{ps}}(\tau)(z_{x}+z_{y}-2F(\tau))^{{\color[rgb]{0,0,0}2}}$ [42]
@@
where $F(\tau)$ may be any time-(in)dependent function. The tilted two-point probability,
where $Z_{xy}(s)=[r_{xy}p_{x}(\tau)-r_{yx}p_{y}(\tau)]/[r_{xy}p_{x}(\tau)+r_{yx}p_{y}(\tau)]$ and $\Sigma_{\mathrm{ps}}=\sum_{x,y}p_{x}(\tau)r_{xy}Z_{xy}(\tau)^{2}$, requires knowledge about transition rates, hence the pseudo variance is not operationally accessible. However, it can be controlled by choosing bounded observables. From the expression Eq. (23), we can see that $p_{xy}^{\mathrm{ps}}(\tau)$ tilts the probability weight onto dissipating transitions. Hence, the pseudo-variance projects the observable onto dissipating transitions. Further, we emphasize that the pseudo-variance is a mathematical construct, which needs to be bounded further, hence the physical interpretation is not relevant. In NESS, Eq. (22) reads
@@
The response of a single-time observable $O(t)$ of a continuous-space diffusion to a perturbation in the noise amplitude, which effectively is a temperature perturbation [91], $\bm{\sigma}\to\bm{\sigma}+\delta\tilde{\bm{\sigma}}$ with perturbation noise amplitude $\tilde{\bm{\sigma}}$ was recently derived using stochastic calculus and reads [86]
@@
Note that similar results have recently been obtained in with a martingale approach [46].
@@
The result in Eq. (29) relates the response to an observable $O$ to a general small perturbation of the dynamics to correlation functions of the unperturbed dynamics (which, notably, for both MJP and diffusions are not necessarily operationally accessible). In the case of equilibrium dynamics, Eq. (29) reduces to the Fluctuation Dissipation Theorem (FDT) in Eq. (33). Therefore, Eq. (29) can in some sense be seen as a (mathematical) extension of the FDT to irreversible and transient systems.
We further generalize the response in Eq. (29) to perturbations of pathwise observables of MJP (see Appendix A.8). Without loss of generality we set
$O_{{\color[rgb]{0,0,0}1}}(t)$ $=\int_{s=0}^{s=t}\mathrm{Tr}[\bm{b}(s)^{T}\mathrm{d}\bm{\varepsilon}(s)],$
${\color[rgb]{0,0,0}O_{2}(t)}$ ${\color[rgb]{0,0,0}=\int_{0}^{t}g_{s}\mathrm{d}s,}$
for some time-dependent hollow matrix $\bm{b}(s)$ and state function $g_{s}\equiv\sum_{x}\delta_{x_{s}x}g_{x}(s)$. For a perturbation strength $h(s)$, i.e., $r_{xy}\to r_{xy}+h(s)\tilde{r}_{xy}$, the result Eq. (29) for $O_{\rm p}(t)=O_{1}(t)+O_{2}(t)$ and its path-wise extension in Sec. V.3.3 can be recast into the familiar response-function formalism
$\Delta O_{\rm p}(t)$ $\equiv\langle O_{\rm p}(t)\rangle_{h}-\langle O_{\rm p}(t)\rangle$
$=\int_{0}^{t}\mathrm{d}sh(s)\chi_{O_{\rm p}}(t,s)+\mathcal{O}(h^{2}),$ (34)
with response function $\chi_{O_{\rm p}}(t,s)$
$\chi_{O_{\rm p}}(t,s)$ $=\chi_{O_{1}}(t,s)+\chi_{O_{2}}(t,s)$ (35)
$={\color[rgb]{0,0,0}\mathbb{1}}_{s<t}\sum_{x,y\neq x}b_{xy}(s)\tilde{r}_{xy}p_{x}(s)+\int_{0}^{t}\mathrm{d}{\color[rgb]{0,0,0}\mathbb{1}}_{z>s}\langle g_{z}R(s)\rangle\,$,
with $R_{x}(s)=\sum_{y}\tilde{r}_{yx}p_{y}(s)/p_{x}(s)$.
@@
Stochastic Calculus and Open Quantum Systems
Open discrete quantum systems coupled to a thermal bath are described by the Belavkin equation [27]
$\mathrm{d}\rho_{s}^{\mathrm{u}}=\mathcal{B}(\rho_{s}^{\mathrm{u}})\mathrm{d}s+\sum_{i}\left(\frac{\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})}{\mathrm{Tr}[\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})]}-\rho_{s}^{\mathrm{u}}\right)\mathrm{d}\tilde{n}_{is},$ (36)
where the pure state $\rho_{s}^{\mathrm{u}}$ is a rank-1 matrix and $\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})=\hat{J}_{i}\rho_{s}^{\mathrm{u}}\hat{J}_{i}^{\dagger}$ with jump operator $\hat{J}_{i}$ for the $i$th quantum jump. The deterministic drift in Eq. (36) has the form
$\mathcal{B}(\rho_{s}^{\mathrm{u}})=-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}-\rho_{s}^{\mathrm{u}}\mathrm{Tr}(-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}),$ (37)
where the the effective Hamiltonian $\hat{H}_{\mathrm{eff}}=\hat{H}-\frac{i}{2}\sum_{j}\hat{J}_{j}^{\dagger}\hat{J}_{j}$ includes the Hamiltonian $\hat{H}$ and $\mathrm{d}\tilde{n}_{is}$ is a stochastic differential which takes values 1 if the $i$th transition occurs in $s,s+\mathrm{d}s]$ and 0 otherwise. The statistical properties of the noise are $\langle\mathrm{d}\tilde{n}_{is}\rangle=\mathrm{Tr}(\mathcal{J}_{i}(\rho_{s}))$ and $\mathrm{d}\tilde{n}_{is}\mathrm{d}\tilde{n}_{js}=\delta_{ij}\mathrm{d}\tilde{n}_{is}$.
We show that the classical analogue to Eq. (36) reads
$\mathrm{d}\delta_{x_{s}x}=$ $\sum_{y\neq x}\left(r_{yx}\delta_{x_{s}y}\mathrm{d}s-\langle\mathrm{d}n_{yx}\rangle-r_{xy}\delta_{x_{s}x}\mathrm{d}s+\langle\mathrm{d}n_{xy}\rangle\right)$
$+\sum_{y\neq x}(\mathrm{d}\varepsilon_{yx}+\langle\mathrm{d}n_{yx}\rangle-\mathrm{d}\varepsilon_{xy}-\langle\mathrm{d}n_{xy}\rangle).$ (38)
Taking the average over all noise histories in Eqs. (36) and (38) recovers the Lindblad and master equation, respectively. Hence, Eqs. (36) and (38) are the stochastic evolution of states in the system. These, as well as other functionals of stochastic increments, we can study as we know the underlying equations of motion, e.g., Eq. (2).
@@ III.1 Stochastic Equations of Motion
Let $(\bm{x}_{\tau})_{0\leq\tau\leq t}$ be a trajectory of $\bm{x}_{\tau}\in{\color[rgb]{0,0,0}\mathbb{R}}^{d}$ evolving according to overdamped Langevin dynamics [51, 144, 86]
@@
Consider now a discrete state space $\mathcal{N}\subseteq{\color[rgb]{0,0,0}\mathbb{N}}$ which, without loss of generality, we can enumerate as $\mathcal{N}=\{1,2,\dots,N\}$. Focusing for the time being on time-homogeneous systems (i.e., time-independent transition rates $r_{ij}$ between states $i$ and $j$) the state probability vector $p_{i}(\tau)$ evolves according to the master equation [140, 145]
@@
where the summation is only over $j$. In terms in the incoming and outgoing fluxes $J_{xy}(\tau)=r_{xy}p_{x}(\tau)$ we have as $\frac{\mathrm{d}}{\mathrm{d}\tau}p_{i}(\tau){\color[rgb]{0,0,0}=}\sum_{j\neq i}[J_{ij}(\tau)-J_{ji}(\tau)]$, and in vector notation $\frac{\mathrm{d}}{\mathrm{d}\tau}\bm{p}(\tau)=\bm{L}\bm{p}(\tau)$ with generator $\bm{L}$ having entries $L_{ij}=\delta_{ij}r_{ii}+(1-\delta_{ij})r_{ji}$ and exit rate $r_{ii}=-\sum_{j\neq i}r_{ij}$. The transition probability (or “propagator”) is therefore given by $P(i,\tau^{\prime}-\tau|j)=P(i,\tau^{\prime}|j,\tau)=\left(\exp{(\tau^{\prime}-\tau)\bm{L}}\right)_{ij}$.
A path is fully determined by the sequence of states $(x_{i})$ and transition times $t_{i}$, $(x_{\tau})_{0\leq\tau\leq t}=(x_{0},t_{0}=0;x_{1},t_{1};\dots)$. The normalized probability of such a trajectory is then [144, 109, 95, 69, 147]
${\color[rgb]{0,0,0}\mathbb{P}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]=p_{x_{0}}(0)\mathrm{e}^{\sum_{k}\tau_{k}(t)r_{kk}}\prod_{i,j}(r_{ij})^{n_{ij}(t)},$ (41)
@@
$\underbrace{\mathrm{d}n_{xy}(\tau)}_{\text{``displacement"}}=\underbrace{r_{xy}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}}_{\text{``drift"}}\mathrm{d}\tau+\underbrace{\mathrm{d}\varepsilon_{xy}(\tau)}_{\text{``noise"}}.$ (46)
Defining the matrix $(\bm{R}(x_{\tau}))_{ij}\equiv(1-\delta_{ij})r_{ij}{\color[rgb]{0,0,0}\delta_{x_{\tau}i}}$, we can write Eq. (46) as a stochastic matrix equation
@@
Meaning Continuous Discrete
Displacement $\mathrm{d}\bm{x}_{\tau}$ $\mathrm{d}\bm{n}(\tau)$
Noise $\bm{\sigma}\mathrm{d}\bm{W}_{\tau}$ $\mathrm{d}\bm{\varepsilon}(\tau)$
Note that $\mathrm{d}\bm{\varepsilon}(\tau)$ is not Gaussian noise, but rather a shifted (due to the $\bm{R}(x_{\tau})\mathrm{d}\tau$ term) time-inhomogeneous Poisson process. Explicitly, because for $x_{\tau}=x$ and $y\neq x$ the differential $\mathrm{d}n_{xy}(\tau)\sim\mathrm{Poi}(r_{xy}\mathrm{d}\tau)$ is Poisson, the values $\mathrm{d}\varepsilon_{xy}(\tau)$ can take for any $x_{\tau}$ are $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau,2-r_{xy}\mathrm{d}\tau,3-r_{xy}\mathrm{d}\tau,\dots\}$, where $\mathrm{d}\varepsilon_{xy}(\tau)=0$ happens if $x_{\tau}\neq x$. In case $x_{\tau}=x$, one can identify $P(\mathrm{d}\varepsilon_{xy}(\tau)=k-r_{xy}\mathrm{d}\tau|x_{\tau}=x)=\mathrm{poi}_{r_{xy}\mathrm{d}\tau}(k)$ for any $k\in{\color[rgb]{0,0,0}\mathbb{N}}_{\geq 0}$, where $\mathrm{poi}_{\lambda}(k)=\lambda^{k}\mathrm{e^{-\lambda}/k!}$ is the Poisson distribution function. In the limit of small $\mathrm{d}\tau$, to linear order in $\mathrm{d}\tau$, the noise therefore takes values $\mathrm{d}\varepsilon_{xy}(\tau)\in\{0,-r_{xy}\mathrm{d}\tau,1-r_{xy}\mathrm{d}\tau\}$ with probabilities $1-p_{x}(\tau),\left(1-r_{xy}\mathrm{d}\tau\right)p_{x}(\tau),r_{xy}\mathrm{d}\tau p_{x}(\tau)$, respectively. The probability for any $k\geq 2$ are $\mathcal{O}(\mathrm{d}\tau^{k})$ and can therefore be neglected.
@@
$=\delta_{xk}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\left[P(i,\tau^{\prime}|l,\tau)-P(i,\tau^{\prime}|k,\tau)\right]r_{kl}+\mathcal{O}(\mathrm{d}\tau).$
@@
The above correlators allow, for example, to evaluate the covariance between the number of jumps at different space-time points along the trajectory, i.e., for $\tau\leq\tau^{\prime}$, [93]
@@
$\times\left[\delta(\tau-\tau^{\prime})\delta_{ik}\delta_{jl}+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau=\tau^{\prime}}\delta_{ik}r_{ij}+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{ij}P(i,\tau^{\prime}|l,\tau)\right].$
@@
${\color[rgb]{0,0,0}\mathbb{P}}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}\mathrm{d}n_{xy}(s)\right],$ (52)